# Adaptive Reduced Tempering For Bayesian inverse problems and rare event simulation

### v1

@@ 3.1 Approximation of a flow of tempered distribution
The sequence of inverse temperatures $\beta_{0}=\beta_{\mathrm{bridge}}(k)<\beta_{1}<\ldots<\beta_{n}=\beta^{(k)}$ will be redefined in this SMC simulation framework as an approximation targeting the sequence defined in (2.13), with the final element being an approximation of the critical inverse temperature (2.14) . However, before going into the latter approximations, let us begin by describing how starting form ${\mu^{N,{(k)}}_{\beta_{i}}}$, SMC simulates at the $i$-th iteration the empirical distribution ${\mu^{N,{(k)}}_{\beta_{i+1}}}$. SMC performs sucessively: i) a weighting and selection step of the $N$ particles, and ii) a mutation step of each individual particle. The weighting step i) assigns weights to the empirical distribution ${\mu^{N,{(k)}}_{\beta_{i}}}$ so that it approximates $\mu_{\beta_{i+1}}^{(k)}$
@@
The above weighting step is followed by a selection step, where the particles are resampled according to their weights $w_{\ell}^{(i)}$. For this purpose, we choose the efficient systematic resampling scheme among the multiple available resampling techniques [18]. The mutation step ii) then consists of a Markovian transition that leaves $\mu_{\beta_{i+1}}^{(k)}$ invariant and is applied individually and independently a certain number of times to the $N$ particles. In practice, the Markovian transitions are performed using a Metropolis-Hasting algorithm, with an adaptive Markovian kernel in order to maintain a given Metropolis acceptance rate [19].
We emphasize that tempering with SMC typically involves many evaluations of the score function underlying the sequence of targeted distributions. This paper precisely focuses on cases where the true score $S^{\ast}$ too expensive for so many evaluation, is substituted by but a reduced score $S^{(k)}$, much cheaper to evaluate.
@@ 3.2 Approximation of critical temperature and adaptive tempering
The simulation starts from an i.i.d. sample with the reference distribution ($\beta_{\mathrm{bridge}}(1)=0$ so that $\mu_{\beta_{\mathrm{bridge}}(1)}^{(1)}=\pi$). We recall that after $k$ iterations, the algorithm provides a current reduced score $S^{(k)}$, an associated pointwise error estimate $E^{(k)}$, an initial inverse temperature $\beta_{\mathrm{bridge}}(k)$ and a sample of size $N$ approximately distributed according to $\mu_{\beta_{\mathrm{bridge}}(k)}^{(k)}$, denoted using a sample-size superscript by $\mu_{\beta_{\mathrm{bridge}}(k)}^{N,(k)}$.
@@
$\mathrm{Ent}\left(\check{\eta}_{\beta}^{(k)}\mid\mu_{\beta}^{(k)}\right)$ $=\log\frac{Z_{\beta}^{(k)}}{\check{Z}_{\beta}^{(k)}}-\beta\check{\eta}_{\beta}^{(k)}(E^{(k)}),$
$\approx\log\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}-\beta E^{(k)}})}-\beta\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}-\beta E^{(k)}}E^{(k)})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}-\beta E^{(k)}})},$ (3.1)
@@
$\mathrm{Ent}\left(\mu_{\beta}^{(k)}\mid\mu_{\beta_{i}}^{(k)}\right)$ $=\log\frac{Z_{\beta_{i}}^{(k)}}{Z_{\beta}^{(k)}}+(\beta-\beta_{i}){\mu}_{\beta}^{(k)}(S^{(k)}),$
$\approx-\log{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}})}{}+(\beta-\beta_{i})\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}}S^{(k)})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}})},$ (3.2)
@@
$\delta\beta=\min\{\delta\beta^{\prime}:\delta\beta^{\prime}\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta^{\prime}S^{(k)}}S^{(k)})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta^{\prime}S^{(k)}})}-\log{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta^{\prime}S^{(k)}})}>c_{2},0\leq\delta\beta^{\prime}\leq\beta_{\infty}-\beta_{i}\}.$
@@
$\log\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta S^{(k)}})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta S^{(k)}-(\beta_{i}+\delta\beta)E^{(k)}})}-(\beta_{i}+\delta\beta)\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta S^{(k)}-(\beta_{i}+\delta\beta)E^{(k)}}E^{(k)})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta S^{(k)}-(\beta_{i}+\delta\beta)E^{(k)}})}\leq c_{1}.$ (3.3)
@@ 3.3 Approximation for bridging
As exposed in Section 2.5, the bridging procedure at the end of iteration $k$ consists in selecting a past index $=k_{\mathrm{bridge}}\leq k$ and a new initial inverse temperature ${\beta_{\mathrm{bridge}}(k+1)}$ which will be used to initialize the SMC simulation with the new score approximation $S^{(k+1)}$ starting from the past, recorded, empirical distribution $\mu^{N,{(k_{\mathrm{bridge}})}}_{\beta^{(k_{\mathrm{bridge}})}}$. In line with the definition of problem (2.8), we select for this purpose the largest lexicographic pair $(k^{\prime},\beta)\in\llbracket 0,k\rrbracket\times\mathcal{C}(k^{\prime})$. However, in order to provide an implementation of this step of the ART algorithm, it is necessary to define an empirical counterpart to this optimization problem. As done previously, using $\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}$, the worst log-cost is approximated for $\beta$ close to ${\beta^{(k^{\prime})}}$ by
$\mathrm{Ent}\left(\check{\eta}_{\beta}^{(k+1)}\mid\mu_{\beta}^{(k+1)}\right)$ $\approx\log\frac{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}})}{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}-\beta E^{(k+1)}})}-\beta\frac{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}-\beta E^{(k+1)}}E^{(k+1)})}{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}-\beta E^{(k+1)}})},$ (3.4)
@@
$\mathrm{Ent}\left(\mu_{\beta}^{(k+1)}\mid\mu_{\beta^{(k^{\prime})}}^{(k^{\prime})}\right)\approx-\log{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}})}+\frac{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}}{\phi_{\beta,k^{\prime}}})}{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}})},$ (3.5)
with
$\phi_{\beta,k^{\prime}}={{\beta S^{(k+1)}-\beta^{(k^{\prime})}S^{(k^{\prime})}}}.$
A SMC approximation of the bridging procedure (2.8) consists then in solving a sequence of one-dimensional constrained optimization problem of the form
@@
We remark that the set of constraints related to this problem is not necessarily convex. However, we can efficiently solved this one-dimensional problem with standard optimization techniques. We choose to use an interior point method. Analogously to the procedure detailed in Section 3.1, the normalization estimate is then updated as
@@ 3.4 SMC-based estimators
Using the sequence of empirical distributions $\{\mu_{{\beta^{(k)}}}^{{N,(k)}}\}_{k=1}^{K}$ generated by SMC simulations along the ART algorithm iterates, we can write a closed-form approximation of the importance sampling estimator given in (2.1) - (2.2). The SMC approximation of the IS estimator $\hat{\gamma}^{H}_{\beta_{\infty}}$ of the the unnormalized distribution $\gamma^{\ast}_{\beta_{\infty}}$ is given by
@@
Moreover, SMC simulation also provides as by-products an estimator of $\gamma^{\ast}_{\beta_{\infty}}$, which we will call reduced SMC (RSMC) estimator because based solely on the reduced score, given by
@@
$\hat{\eta}^{N,H}_{\beta_{\infty},IS}=\hat{\gamma}^{N,H}_{\beta_{\infty},IS}/\hat{\gamma}^{N,H}_{\beta_{\infty},IS}(\,{\rm l}{\rm l})\quad\textrm{and}\quad\hat{\eta}^{N,H}_{\beta_{\infty},RSMC}=\hat{\gamma}^{N,H}_{\beta_{\infty},RSMC}/\hat{\gamma}^{N,H}_{\beta_{\infty},RSMC}(\,{\rm l}{\rm l}).$ (3.9)
@@ 4 Numerical Simulations
Our numerical simulations focus on target distributions of the form (1.1), where $S^{\ast}$ implies the high-dimensional solution of an elliptic PDE parametrized by some random input vector. We propose to evaluate numerically the ART algorithm for sampling these distributions, where the PDE solution is approximated by a popular reduced basis (RB) approach. The evaluation is completed by a study of the ART’s performance in the context of a one-dimensional yet challenging toy problem. Our numerical simulations are carried out using an implementation of the ART algorithm in the Julia language, available at https://gitlab.inria.fr/pheas/adaptive-reduced-tempering. After introducing the models, namely the PDE and it high-fidelity and reduced solutions, we will study two related problems: a rare event estimation problem and a Bayesian inverse problem.
@@ 4.1 Model #RB: heat diffusion and reduced basis approximations
Figure 1: Illustration of the thermal block problem (Model #RB). High-fidelity solution $\Psi^{\ast}(X)$ of the PDE for $4$ different samples $X$ of the reference log-normal distribution (images taken from [11]).
Let $\Psi^{\ast}(x)$ be the high-fidelity numerical approximation of the solution of a bi-dimensional PDE on the spatial domain $\Omega=[0,1]^{2}$ parametrized by a $4$-dimensional random vector $x=(x_{1}\cdots x_{4})^{\intercal}\in\mathcal{X}=\mathbb{R}_{+}^{4}$. The chosen PDE corresponds to the well-studied thermal block problem, which models heat diffusion in an heterogeneous media.
@@
$-\nabla\cdot(\kappa(\cdot;x)\nabla\Psi)=&1\quad\textrm{in}\quad\Omega\\ \Psi=&0\quad\textrm{on}\quad\partial\Omega\end{aligned}\right.,$
@@
The high-fidelity solution $\Psi^{\ast}(x)$ of this diffusion problem entails the solution of a variational problem. Precisely, let $\mathcal{V}$ denote the Hilbert space $\mathcal{H}^{1}_{0}(\Omega)$ and $\|\nu\|_{\mathcal{V}}=\|\nabla\nu\|_{L^{2}}$ the norm induced by the inner product defined over $\mathcal{V}$. For any $x\in\mathcal{X}$, we define the solution $\Psi^{\star}(x)\in\mathcal{V}$ of the variational problem $a(\Psi(x),\nu;x)=\int_{\Omega}\nu\mathrm{d}\Omega,\quad\forall\nu\in\mathcal{V},$ where $a:\mathcal{V}\times\mathcal{V}\to\mathbb{R}$ is the bilinear form $a(u,\nu;x)=\int_{\Omega}\kappa(\cdot;x)\nabla u\cdot\nabla\nu d\Omega$. We consider the piecewise constant diffusion coefficient $\kappa(\cdot;x)=\sum_{q=1}^{4}\,{\rm l}{\rm l}_{\Omega_{q}}(\cdot)x_{q}$ with the partition $\Omega=\cup_{q=1}^{4}\Omega_{q}$ and $\Omega_{q}$’s being non-overlapping squares. Setting the parameter space $\mathcal{X}=(\mathbb{R}_{+})^{4}$ implies that the bilinear form is strongly coercive, and the Lax-Milgram lemma ensures the variational problem has a unique solution.
Consider next a finite $h$-dimensional subspace $\mathcal{V}_{h}\subset\mathcal{V}$. The high-fidelity approximation $\Psi^{\ast}(x)\in\mathcal{V}_{h}$ of the solution $\Psi^{\star}(x)$ is defined for any $x\in\mathcal{X}$ by
@@
with the energy norm induced by the bilinear form $\|\nu\|=\sqrt{a(\nu,\nu;x)}$. The high-fidelity approximation $\Psi^{\ast}(x)$ is obtained in practice by solving a large system of $h$ linear equations. [29] Figure 1 shows several high-fidelity solutions $\Psi^{\ast}(x)$ of the PDE for samples drawn according to the reference distribution.
@@
=$\mathcal{M}=\{\Psi^{\ast}(x)\in\mathcal{V}_{h}:x\in\mathcal{X}\},$
@@
The computation of the RB approximation and error bound relies on the affine parametric dependence of the RB approximation by splitting the assembly of the reduced matrices and vectors in an off-line and on-line phases. The former, which is performed once for all at each RB update, entails the computation of all the $h$-dependent and $x$-independent structures based on the snapshots [^4] With regard to the snapshots used to enrich the RB approximation space, it should be noted that the sampling of snapshots in the first iterations of ART corresponds to the popular weak greedy algorithm. However, once the target temperature has been reached ($j_{0}$ times), the weak greedy algorithm is replaced by drawing snapshots according to the importance distribution. . It requires $\mathcal{O}(Kh^{2}+K^{2}h)$ operations. Besides, one must add the evaluation of the snapshot, which is $\mathcal{O}(h^{3})$. In the latter, on-line phase, for a given parameter $x\in\mathcal{X}$, we assemble and solve the RB system and provide the error bound in $\mathcal{O}(K^{3})$, with a cost depending only on $K$. For further details, we refer the reader to the introductory book on RB [29].
For sampling problems involving this PDE model, the reference distribution $\pi$ for $x$ will be a log-normal distribution of independent components with $\ln(x_{q})\sim\mathcal{N}(0.6,0.8^{2})$, for $q=1,\ldots,4$.

### v2

@@ 3.1 Approximation of a flow of tempered distribution
The sequence of inverse temperatures $\beta_{0}=\beta_{\mathrm{bridge}}(k)<\beta_{1}<\ldots<\beta_{n}=\beta^{(k)}$ will be redefined in this SMC simulation framework as an approximation targeting the sequence defined in (2.13), with the final element being an approximation of the critical inverse temperature (2.14). However, before going into the latter approximations, let us begin by describing how starting from $\mu^{N,{(k)}}_{\beta_{i}}$}, SMC simulates at the $i$-th iteration the empirical distribution $\mu^{N,{(k)}}_{\beta_{i+1}}$}. SMC performs successively: i) a weighting and selection step of the $N$ particles, and ii) a mutation step of each individual particle. The weighting step i) assigns weights to the empirical distribution $\mu^{N,{(k)}}_{\beta_{i}}$} so that it approximates $\mu_{\beta_{i+1}}^{(k)}$
@@
The above weighting step is followed by a selection step, where the particles are resampled according to their weights $w_{\ell}^{(i)}$. For this purpose, we choose the efficient systematic resampling scheme among the multiple available resampling techniques [21]. The mutation step ii) then consists of a Markovian transition that leaves $\mu_{\beta_{i+1}}^{(k)}$ invariant and is applied individually and independently a certain number of times to the $N$ particles. In practice, the Markovian transitions are performed using a Metropolis-Hastings algorithm, with an adaptive Markovian kernel in order to maintain a given Metropolis acceptance rate [23].
We emphasize that tempering with SMC typically involves many evaluations of the score function underlying the sequence of targeted distributions. This paper precisely focuses on cases where the true score $S^{\ast}$ is too expensive for so many evaluations and is substituted by a reduced score $S^{(k)}$, much cheaper to evaluate.
@@ 3.2 Approximation of critical temperature and adaptive tempering
The simulation starts from an i.i.d. sample with the reference distribution ($\beta_{\mathrm{bridge}}(1)=0$ so that $\mu_{\beta_{\mathrm{bridge}}(1)}^{(1)}=\pi$). We recall that after $k$ iterations, the algorithm provides a current reduced score $S^{(k)}$, an associated point-wise error estimate $E^{(k)}$, an initial inverse temperature $\beta_{\mathrm{bridge}}(k)$ and a sample of size $N$ approximately distributed according to $\mu_{\beta_{\mathrm{bridge}}(k)}^{(k)}$, denoted using a sample-size superscript by $\mu_{\beta_{\mathrm{bridge}}(k)}^{N,(k)}$.
@@
$\mathrm{Ent}\left(\check{\eta}_{\beta}^{(k)}\mid\mu_{\beta}^{(k)}\right)$ $=\ln\frac{Z_{\beta}^{(k)}}{\check{Z}_{\beta}^{(k)}}-\beta\check{\eta}_{\beta}^{(k)}(E^{(k)}),$
$\approx\ln\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}-\beta E^{(k)}})}-\beta\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}-\beta E^{(k)}}E^{(k)})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}-\beta E^{(k)}})},$ (3.1)
@@
$\mathrm{Ent}\left(\mu_{\beta}^{(k)}\mid\mu_{\beta_{i}}^{(k)}\right)$ $=\ln\frac{Z_{\beta_{i}}^{(k)}}{Z_{\beta}^{(k)}}+(\beta-\beta_{i}){\mu}_{\beta}^{(k)}(S^{(k)}),$
$\approx-\ln{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}})}{}+(\beta-\beta_{i})\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}}S^{(k)})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{(\beta-\beta_{i})S^{(k)}})},$ (3.2)
@@
$\delta\beta=\min\{\delta\beta^{\prime}:\delta\beta^{\prime}\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta^{\prime}S^{(k)}}S^{(k)})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta^{\prime}S^{(k)}})}-\ln{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta^{\prime}S^{(k)}})}>c_{2},0\leq\delta\beta^{\prime}\leq\beta_{\infty}-\beta_{i}\}.$
@@
$\ln\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta S^{(k)}})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta S^{(k)}-(\beta_{i}+\delta\beta)E^{(k)}})}-(\beta_{i}+\delta\beta)\frac{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta S^{(k)}-(\beta_{i}+\delta\beta)E^{(k)}}E^{(k)})}{\mu^{N,{(k)}}_{\beta_{i}}(e^{\delta\beta S^{(k)}-(\beta_{i}+\delta\beta)E^{(k)}})}\leq c_{1}.$ (3.3)
@@ 3.3 Approximation for bridging
As exposed in Section 2.5, the bridging procedure at the end of iteration $k$ consists in selecting a past index $k_{\mathrm{bridge}}\leq k$ and a new initial inverse temperature $\beta_{\mathrm{bridge}}(k+1)$} which will be used to initialize the SMC simulation with the new score approximation $S^{(k+1)}$ starting from the past, recorded, empirical distribution $\mu^{N,{(k_{\mathrm{bridge}})}}_{\beta^{(k_{\mathrm{bridge}})}}$. In line with the definition of problem (2.8), we select for this purpose the largest lexicographic pair $(k^{\prime},\beta)\in\llbracket 0,k\rrbracket\times\mathcal{C}(k^{\prime})$. However, in order to provide an implementation of this step of the ART algorithm, it is necessary to define an empirical counterpart to this optimization problem. As done previously, using $\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}$, the worst log-cost is approximated for $\beta$ close to $\beta^{(k^{\prime})}$} by
$\mathrm{Ent}\left(\check{\eta}_{\beta}^{(k+1)}\mid\mu_{\beta}^{(k+1)}\right)$ $\approx\ln\frac{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}})}{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}-\beta E^{(k+1)}})}-\beta\frac{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}-\beta E^{(k+1)}}E^{(k+1)})}{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}-\beta E^{(k+1)}})},$ (3.4)
@@
$\mathrm{Ent}\left(\mu_{\beta}^{(k+1)}\mid\mu_{\beta^{(k^{\prime})}}^{(k^{\prime})}\right)\approx-\ln{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}})}+\frac{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}}{\phi_{\beta,k^{\prime}}})}{\mu^{N,{(k^{\prime})}}_{\beta^{(k^{\prime})}}(e^{\phi_{\beta,k^{\prime}}})},$ (3.5)
with $\phi_{\beta,k^{\prime}}={{\beta S^{(k+1)}-\beta^{(k^{\prime})}S^{(k^{\prime})}}}.$ A SMC approximation of the bridging procedure (2.8) consists then in solving a sequence of one-dimensional constrained optimization problem of the form
@@
We remark that the set of constraints related to this problem is not necessarily convex. However, we can efficiently solve this one-dimensional problem with standard optimization techniques. We choose to use an interior point method. Analogously to the procedure detailed in Section 3.1, the normalization estimate is then updated as
@@ 3.4 SMC-based estimators
Using the sequence of empirical distributions $\{\mu_{{\beta^{(k)}}}^{{N,(k)}}\}_{k=1}^{K}$ generated by SMC simulations along the ART algorithm iterations, we can write a closed-form approximation of the importance sampling estimator given in (2.1) - (2.2). The SMC approximation of the IS estimator $\hat{\gamma}^{H}_{\beta_{\infty}}$ of the the unnormalized distribution $\gamma^{\ast}_{\beta_{\infty}}$ is given by
@@
Moreover, SMC simulation also provides as by-products an estimator of $\gamma^{\ast}_{\beta_{\infty}}$, which we will call reduced SMC (RSMC) estimator because it is based solely on the reduced score, given by
@@
$\hat{\eta}^{N,H}_{\beta_{\infty},IS}=\hat{\gamma}^{N,H}_{\beta_{\infty},IS}/\hat{\gamma}^{N,H}_{\beta_{\infty},IS}(\,{\rm l}\hskip-1.99168pt{\rm l})\quad\textrm{and}\quad\hat{\eta}^{N,H}_{\beta_{\infty},RSMC}=\hat{\gamma}^{N,H}_{\beta_{\infty},RSMC}/\hat{\gamma}^{N,H}_{\beta_{\infty},RSMC}(\,{\rm l}\hskip-1.99168pt{\rm l}).$ (3.9)
@@ 4 Numerical Simulations
Our numerical simulations focus on target distributions of the form (1.1), where $S^{\ast}$ implies the high-dimensional solution of an elliptic PDE parametrized by some random input vector. We propose to evaluate numerically the ART algorithm for the sampling of these distributions, where the PDE solution is approximated by a popular reduced basis (RB) approach. The evaluation is completed by a study of the ART’s performance in the context of a one-dimensional yet challenging toy problem. Our numerical simulations are carried out using an implementation of the ART algorithm in the Julia language, available at https://gitlab.inria.fr/pheas/adaptive-reduced-tempering. After introducing the models, namely the PDE and its high-fidelity and reduced solutions, we will study two related problems: a rare event estimation problem and a Bayesian inverse problem.
@@ 4.1 Model #RB: heat diffusion and reduced basis approximations
Fig. 2: Illustration of the thermal block problem (Model #RB). High-fidelity solution $\Psi^{\ast}(X)$ of the PDE with $d=4$ for 44 different samples $X$ of the reference log-normal distribution (images taken from [13]).
Let $\Psi^{\ast}(x)$ be the high-fidelity numerical approximation of the solution of a bi-dimensional PDE on the spatial domain $\Omega=[0,1]^{2}$ parametrized by a $d$-dimensional random vector $x=(x_{1}\cdots x_{d})^{\intercal}\in\mathcal{X}=\mathbb{R}_{+}^{d}$. We will consider $d=2\times 2$ or $d=5\times 5$ in our experiments. The chosen PDE corresponds to the well-studied thermal block problem, which models heat diffusion in a heterogeneous medium.
@@
$\left\{\begin{aligned} -\nabla\cdot(\kappa(\cdot;x)\nabla\Psi)=&1\quad\textrm{in}\quad\Omega\\ \Psi=&0\quad\textrm{on}\quad\partial\Omega\end{aligned}\right$.,
@@
The high-fidelity solution $\Psi^{\ast}(x)$ of this diffusion problem entails the solution of a variational problem. Precisely, let $\mathcal{V}$ denote the Hilbert space $\mathcal{H}^{1}_{0}(\Omega)$ and $\|\nu\|_{\mathcal{V}}=\|\nabla\nu\|_{L^{2}}$ the norm induced by the inner product defined over $\mathcal{V}$. For any $x\in\mathcal{X}$, we define the solution $\Psi^{\star}(x)\in\mathcal{V}$ of the variational problem $a(\Psi(x),\nu;x)=\int_{\Omega}\nu\mathrm{d}\Omega,\quad\forall\nu\in\mathcal{V},$ where $a:\mathcal{V}\times\mathcal{V}\to\mathbb{R}$ is the bilinear form $a(u,\nu;x)=\int_{\Omega}\kappa(\cdot;x)\nabla u\cdot\nabla\nu d\Omega$. We consider the piecewise constant diffusion coefficient $\kappa(\cdot;x)=\sum_{q=1}^{d}\,{\rm l}\hskip-1.99168pt{\rm l}_{\Omega_{q}}(\cdot)x_{q}$ with the partition $\Omega=\cup_{q=1}^{d}\Omega_{q}$ and $\Omega_{q}$’s being non-overlapping squares. Setting the parameter space $\mathcal{X}=(\mathbb{R}_{+})^{d}$ implies that the bilinear form is strongly coercive, and the Lax-Milgram lemma ensures the variational problem has a unique solution.
Consider next a finite $h$-dimensional subspace $\mathcal{V}_{h}\subset\mathcal{V}$. We will set $h=5101$ in our experiments. The high-fidelity approximation $\Psi^{\ast}(x)\in\mathcal{V}_{h}$ of the solution $\Psi^{\star}(x)$ is defined for any $x\in\mathcal{X}$ by
@@
with the energy norm induced by the bilinear form $\|\nu\|=\sqrt{a(\nu,\nu;x)}$. The high-fidelity approximation $\Psi^{\ast}(x)$ is obtained in practice by solving a large system of $h$ linear equations [38]. Figure 2 shows several high-fidelity solutions $\Psi^{\ast}(x)$ of the PDE for samples drawn according to the reference distribution.
@@
$\mathcal{M}=\{\Psi^{\ast}(x)\in\mathcal{V}_{h}:x\in\mathcal{X}\},$
@@
The on-line computation of the RB approximation and error bound relies on the affine parametric dependence of the RB approximation and the off-line assembly of the reduced matrices and vectors. This off-line phase, which is performed once for all at each RB update, entails the computation of all the $h$-dependent and $x$-independent structures based on the snapshots [^5] With regard to the snapshots used to enrich the RB approximation space, it should be noted that the sampling of snapshots in the first iterations of ART corresponds to the popular weak greedy algorithm. However, once the target temperature has been reached ($j_{0}$ times), the weak greedy algorithm is replaced by drawing snapshots according to the importance distribution. . We consider the case of a Galerkin RB approximation with $K\ll h$ and $d\ll h$. In the case of the finite element discretization of our elliptic problem, computing scalar products, matrix-vector products or the high-fidelity solution has a complexity of $O(h)$, but with a very large constant (thanks to the generalized minimal residual method [1] used to solve sparse linear system in pymor). It thus requires off-line $\mathcal{O}(hdK^{2})$ operations for computing reduced matrices and vectors used to assemble the RB system, and $\mathcal{O}(hd^{2}K^{2})$ operations for computing those used to build the error estimator. Computing the snapshot in $\mathcal{O}(h)$ operations is in general less expensive in comparison. In the latter, on-line phase, for a given parameter $x\in\mathcal{X}$, we solve the RB system in $\mathcal{O}(K^{3}+dK^{2})$ and provide the error bound in $\mathcal{O}(d^{2}K^{2})$, with a cost independent on $h$. For further details, we refer the reader to the introductory book on RB [38]. We will show in Appendix B that the dependence in $d^{2}$ in the complexities for error update and evaluation makes the ART method non-competitive for $d=25$.
For sampling problems involving this PDE model, the reference distribution $\pi$ for $x$ will be a log-normal distribution of independent components with $\ln(x_{q})\sim\mathcal{N}(0.6,0.8^{2})$, for $q=1,\ldots,d$.